A circular garden with a circumference of 80 feet is surrounded by a walkway by a patio that is 6 feet wide around. Find the radius of the garden and the walkway combined. Fine the area of the walkway alone.
Answer by richwmiller(17219)

2*pi*r=80
r=40/pi=12.7324
12.734+6=18.7324 ft radius of garden and walkway
pi*r^2=area
1102.39 area of garden and walkway
pi*18.7324^2-pi*12.7324=593.098 area of walkway alone